Determine the coordinates of the x-intercepts for the quadratic equation 5x2 – 21x + 4 = 0.
Answer by ewatrrr(24785)


Hi
finding the coordinates of the x-intercepts
5x^2 – 21x + 4 = 0
factoring
(5x -1)(x-4)= 0
(5x -1)=0
x = (1/5)
(x-4)= 0
x = 4
coordinates of the x-intercepts are(.2,0) and (4,0)